Porno for Pyros

(features Perry Farrel on vocals, previously of Jane’s Addiction)

Sadness (2:33)

i got the devil in me it’s just a cloud it’s sadness it’s just a cloud then I find compassion and I find it vanishes these thoughts I have they seem to pass me by like a cloud

i was praying it’s a cloud hope it’s just another name for a prayer it’s just a cloud

i want to find some laughter when you laugh they can’t kill you these thoughts i have they seem to pass me by like a cloud

fuel and release! happiness is good for an hour fuel and release! happiness is good for an hour

i hate him for all the things he’s done! because i love her so! i got the devil in me! it’s sadness! it’s just a cloud

Porno for Pyros (3:06)

came home last night there was fire and smoke on the t.v. cops and the army people running out in the street looting

i took off my clothes and came four times– could not leave myself alone

it was porno for pyros!

got in my car saw the fire and smoke, headed to it just for this night i believe it is good to destroy it!

get back in the car! let’s drive on by everybody bump and grind...

it was porno for pyros

Meija (3:13)

where are you going meija?

you got your money in your hand and you mind already out of the door get her started there ain’t no stopping her you gotta wait till she runs out of money or maybe get to sleep she knows everybody baby took the car out for a drive didn’t come back for nearly two days where’s she been? why’d she go? why’s she back now? we all know! we know why we know why she’s back again we know where’s she goin”’ someplace better than here...

everybody count your money! everybody count your money! everybody count your money!

Cursed Female (3:24)

cursed to be born beautifil, poor and female there’s none that suffer more

she’s the queen of the bees that buzz in the bars fresh as a strawberry and she always loved dancing

cursed to be born beautifil, poor and female there’s none that suffer more

2 girls in 2 nights got caught in the back of the alleyway fresh as strawberries 2 girls in 2 nights when no one came out zipped up their clothes and walked back on home cursed to be born beautifil, poor and female there’s none that suffer more

Cursed Male (3:50)

all the guys that really have the money are too old to have a good time with it

too old to turn the women on too old to have a fast car and drive it

all the guys that really have the money are too old to party or get dirty

too old to turn the women on too old to have a fast car and drive it

Pets (3:36)

children are innocent a teenager’s fucked up in the head adults are even more fucked up and elderlies are like children will there be another race to come along and take over for us? maybe martians could do better than we’ve done we’ll make great pets!

my friend says we’re like the dinosaurs only we are doing ourselves in much faster than they ever did we’ll make great pets!

Bad Shit (2:58)

got some bad shit then i walked the beat in venice

i went outside i been in too long outside for some inspiration did the G-O did the No-Go there were stars there were flies even the flies dropped like... what do flies drop like? stars but i didn’t go down

i went outside i thought i was a chinese blues player outside been in too long did the G-O did the No-Go there were stars there were flies even the flies dropped like... what do flies drop like? stars

but i didn’t go down

Packin’ .25 (4:08)

yesterday morning i went out for a cup of coffee i shaved then i combed my hair a man who didn’t know me said something to my back i stopped to turn around and face it

that is why i pack my .25 where nobody knows right above my boot it’s the law no one there to serve you why not be the hero? why not be your own?

swear i’ll kill you! swear that i’ll kill you the law–it’s the law every man out on the street knows swear that i’ll kill you

i dreamt all yesterday how i might make a man feel with a gun up to his face he’d show respect to me ”i don’t care what you’re thinking i’ll wipe that thought away”

Black Girlfriend (4:33)

ever since the riots all i really wanted was a black girlfriend they don’t play around they’re hard enough to keep any man in line thinking of my pale white skin thinking of her dark and smooth she against me my black girlfriend saw her on the corner where she lived, i asked her ”can you braid my hair?” she and her girlfriends laughed at me, said that ”it was easy but it’ll cost you some” looking out her window it’s so exciting and foreign but i’m staying with my black girlfriend

drivin’ thru the hood in my chevy nova ’62 my arm around my little black girl people on the corner looking in my car wanna do me but i won’t give her back my little black girl ”do you wanna come on in? do you wanna eat some? meet my family?” my black girlfriend

Blood Rag (3:29)

somebody handed me a bloody rag oh that rag got thick and red and i wish and i hope and i hope and i pray but oh! no way no way

how many times can you walk away? so many times–he did it again? i wished and i hoped and i hoped and prayed but oh! no way

Orgasm (4:27)

met a girl who’d never had an orgasm i couldn’t believe she’d never had one i said, ”sit back and get yourself relaxed”

she thought that i was just the one she said, ”i know that you could do it” she had to know got curious what it was like enough to trust me you look so good so good... you look so good to me so good to me

(credits)

Perry Farrell: vocals Peter DiStefano: guitars Martyn Le Noble: bass Stephen Perkins: drums, percussion

Sound additives: Skatemaster Tate and Matt Hyde

All songs written and performed by Porno for Pyros

(c) 1992 I’ll Hit You Back Music/Virgin-EMI Songs BMI Lyrics Reprinted by Permission. All Rights Reserved. Produced by Perry Farrell and Matt Hyde Engineered by Matt Hyde and Rob Seifert Assistant engineer: Damien Wagner Recorded and mixed at Crystal Studios, Los Angeles during the summer of ’92 Mastered at Bernie Grundman’s by Chris Belman Cover Artwork: Perry Farrell and Bill Hofstadter Photography: Bill Hoftstadter Art direction and design: Tom Recchion Lettering: Barrie Goshko Cover Computer Imaging: Cecil Juanarena, Insight Communications

Management: Ted Gardner for Larrakin Management Booking Agent: Don Muller at William Morris, L.A. Publicity: Ted Mico at Shattered, L.A.

Thanks to: Stuart Ross, Nikki Brown, Roger Leonard, Steven Baker, Paul V., Eric Greenspan, the Perkins family, G.K., Jr and Sr, Krissi and Ronnie, D.W., Sabian, Remo, Pro-Mark, Ludwig, Todd Trent, Kiva, Toi Russell, the DiStefano family and friends, Melinda Mara, Nannie and Bram, Steve Thomson, Madea, Casey Niccoli, Bo Gardner, Bill Vuylsteke, Kim and Donovan, Andrew Berliner, Michele Matheson, Mike Cassel and Greg, Danny Laughlin and the Venice boys

Skatemaster Tate appears courtesy of U.K. Acid Jazz

(c) 1993 Warner Bros. Records Inc. Made in U.S.A.

† TEX Version by Victor Moral ( U de Chile ) e-mail:vmoral@araucaria.cec.uchile.cl †